Consider the following measurements of ΔABC.

α = 49.19°,    a = 5.04,    b = 6.17

Solve ΔABC. (Round answers to two decimal places. If there is no solution, enter NO SOLUTION.)

smaller c:

c=

β=

γ=  °

larger c:

c=

β=

γ=
Answer #1

© from given data uc - soft Pier 36 98 always Pier is straight ? B SO B=988 Total angle = 188 Sint - SinB = Sinc ..c=60° Sin(- Total angle = 180° -Af3+= 190 C = (8-49.14- 67.96291 180 - 112.10 A+B+(= 180 C = 180 -- 112:1-4919-1971 C = 6291 (large) Si